OpenStudy (anonymous):

Im trying to write the slope intercept form equation for x+y=3 and give the y intercept

OpenStudy (anonymous):

slope-intercept form is y = mx + b so just transpose x to the other side to isolate y

OpenStudy (anonymous):

y=x+-3

OpenStudy (anonymous):

not exactly. it should be x that's negative not 3 so you have y = -x + 3

OpenStudy (anonymous):

so the 3 is the y intercept then

OpenStudy (anonymous):

right

OpenStudy (phi):

yes , 3 is the y-intercept when you start with x+y=3 and you want the y by itself, add -x to both sides: -x + x + y = -x +3 the -x + x is 0 (you have 1 x and take away 1 x) you get y= -x + 3

OpenStudy (anonymous):

Thank you